OpenStudy (anonymous):

A cylindrical vase has a diameter of 4 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 8 inches. What is the volume of water in the vase?

OpenStudy (anonymous):

26π in3 16π in3 24π in3 32π in3

OpenStudy (anonymous):

@DoomDude @jorge3o5 @Mertsj

OpenStudy (mertsj):

Volume of water and marbles is pi r^2h=pi(2)^2(8)=32pi in^3

OpenStudy (mertsj):

Each marble is a sphere of radius 1. The volume of a sphere is 4/3pir^3 so 6 of the marbles has volume 6(4/3)pi(1)^3=8pi

OpenStudy (mertsj):

If we want the volume of the water, we should take the volume of the marbles and water together and subtract the volume of the marbles. I will leave that to you.

OpenStudy (anonymous):

thanks i got it @Mertsj